# Frontier AI’s Impact on the Cybersecurity Landscape

### v2

@@ Abstract
As frontier AI advances rapidly, understanding its impact on cybersecurity and inherent risks is essential to ensuring safe AI evolution (e.g., guiding risk mitigation and informing policymakers). While some studies review AI applications in cybersecurity, none of them comprehensively discuss AI’s future impacts or provide concrete recommendations for navigating its safe and secure usage. This paper presents an in-depth analysis of frontier AI’s impact on cybersecurity and establishes a systematic framework for risk assessment and mitigation. To this end, we first define and categorize the marginal risks of frontier AI in cybersecurity and then systemically analyze the current and future impacts of frontier AI in cybersecurity, qualitatively and quantitatively. We also discuss why frontier AI likely benefits attackers more than defenders in the short term from equivalence classes, asymmetry, and economic impact. Next, we explore frontier AI’s impact on future software system development, including enabling complex hybrid systems while introducing new risks. Based on our findings, we provide security recommendations, including constructing fine-grained benchmarks for risk assessment, designing AI agents for defenses, building security mechanisms and provable defenses for hybrid systems, enhancing pre-deployment security testing and transparency, and strengthening defenses for users. Finally, we present long-term research questions essential for understanding AI’s future impacts and unleashing its defensive capabilities.
11footnotetext: These authors contributed equally to this work.
@@ 1 Introduction
Frontier artificial intelligence (AI), including foundation models and emerging AI agents, has demonstrated remarkable success across domains such as text generation [2], code generation and debugging [113, 236], scientific discovery [234, 149], and healthcare [44]. Recent works [115, 85, 312] further explore AI agents to accomplish complex tasks beyond content generation, such as automatically resolving software issues [312]. While “frontier AI” has been used to refer to foundation models only [11], we use it to represent broader advanced AI systems, including foundation models like large language models (LLMs) [2, 276, 272, 15] and multi-modal models [293], as well as AI agents [85]. [^1] We use both AI agents and hybrid systems to refer to systems that combine foundation models with non-ML symbolic software components, where foundation models interact with other components.
@@
Motivated by their strong capabilities, the security community has also begun utilizing frontier AI, such as using LLM to facilitate fuzzing [196] and generate patching [332]. Some works further explored building foundation models for security data, such as networking traffic, binary code, and blockchain transactions, achieving better performance than traditional ML in threat detection [117, 98, 320, 220]. Recently, AI-driven cyber reasoning systems from the DARPA AIxCC competition [71] found and patched real-world zero-day vulnerabilities [119]. While frontier AI offers great promise, it also introduces significant risks. First, attackers can leverage LLMs to synthesize malicious code [151, 286] or to launch attacks against systems and humans [90, 271]. Second, AI models themselves have serious security concerns [287, 26], including backdoors [205], jailbreaking [52, 53], prompt injection [181], and data leakage [41].
@@
Figure 1: Performance evolvement of frontier AI models on coding, math, and science benchmarks over time. The coding-related tasks are emphasized in bold text.
Given these successes and concerns, a key question arises: Will frontier AI benefit attackers or defenders more, and how might it reshape the cybersecurity landscape? While recent work surveys LLMs’ applications in cybersecurity [93, 124, 102, 318, 239, 331, 337, 136, 253], they do not provide a comprehensive analysis of how frontier AI will impact the cybersecurity landscape. These works also do not analyze potential risks introduced by frontier AI or provide actionable recommendations on how the security community should react to such risks.
With the rapid AI development, it is critical to analyze frontier AI’s security impact and provide recommendations [1, 249]. Figure 1 shows frontier AI models’ capabilities improving at an unprecedented pace, particularly in code reasoning, generation, and patching. The o3 model—launched just two months after o1—significantly outperformed o1 on Codeforces (2727 vs. 1891) and the SWE-bench Verified benchmark (71.7% vs. 48.9%) [278]. This trend implies that AI could become increasingly leverageable by attackers than previously expected.
In this paper, we address the gap by presenting a comprehensive study of frontier AI’s current and future impact on cybersecurity, as well as a call to action for system builders, defenders, and AI developers. We begin by defining and analyzing the marginal risks [153] of frontier AI to cybersecurity and provide a detailed risk categorization based on the standard cyber kill chain [86]. Using this categorization, we summarize frontier AI’s current use and capability in attacks and defenses and discuss its potential future impacts, quantitatively and qualitatively. Next, we analyze whether frontier AI will benefit attackers or defenders more from the perspectives of equivalence classes, asymmetries between attackers and defenders, and economic impacts. We then explore how frontier AI will reshape future system design. Based on these projections, we offer actionable recommendations for leveraging frontier AI to build more secure systems while assessing and preparing for its potential risks. Finally, we present a set of open research questions about frontier AI and security.
Findings.
  • Our systematic review shows that most AI-driven attacks on software systems are demonstrated in research settings and are not yet widely applied in practice. Conversely, many real-world AI-driven attacks on humans (e.g., phishing, social engineering, deepfakes) are already occurring on a large scale.
  • Frontier AI advances automation in defense, particularly in threat detection and Security Operations Center (SOC) operations, but has not been widely used in vulnerability triage and remediation. Overall, its influence on defense is less than on offense.
  • We argue that frontier AI will likely benefit attackers more than defenders in the short term. However, with a coordinated effort—through enhanced risk assessment, smarter defense integration, secure-by-design system development—the balance could eventually tip in favor of defenders.
  • We expect more hybrid systems combining AI with traditional symbolic components, which will introduce new security challenges.
Call to action.
  • We recommend comprehensive risk assessments and continuous monitoring of frontier AI capabilities and resulting cybersecurity risks. This involves building comprehensive, fine-grained benchmarks, dynamically updating them, and continuously conducting marginal risk assessments.
  • We recommend strategically leveraging frontier AI to strengthen every phase of the cybersecurity defense lifecycle. To this end, we can build defenses combining AI with existing program analysis-based approaches. We can also explore hybrid systems with multiple AI agents to enhance defenses.
  • We recommend more efforts to foster better security for hybrid systems, whether through secure-by-design approaches or by creating specialized proactive/reactive defenses.
@@
  • We recommend AI developers conduct more comprehensive pre-deployment security testing and enforce greater transparency in AI development. Moreover, giving defenders early and differential access can help shift the balance in favor of the defenders over the attackers.
  • To mitigate human-targeted attacks, we call for better education about AI-enabled attacks and security best practices, while advancing research on defenses against sophisticated AI-powered threats targeting humans.
These recommendations aim to guide the community to understand and mitigate frontier AI’s marginal risks, turning it into a powerful tool for strengthening cybersecurity.
@@ 2 Methodology
2.1 Marginal Risk Assessment Framework
Based on the concept of marginal risk [153], which refers to new and unique risks introduced by emerging technologies, we define the marginal risks of frontier AI in cybersecurity as novel cyber threats either introduced or amplified by frontier AI through intentional misuse beyond traditional attack methods.
First, we study how frontier AI amplifies existing attacks against systems and users. For attacks against systems, we follow the MITRE ATT&CK framework [201] and cyber kill chain [190] (Figure 6) to categorize existing attacks into seven steps. We follow the cyber kill chain to name each step and merge the corresponding steps in MITRE ATT&CK into the sub-steps (Figure 2). Similarly, we classify attacks against humans based on their goals and methods (Figure 3). Based on this categorization, we conduct a detailed analysis of the marginal risks of frontier AI to each attack step through qualitative and quantitative analysis.
Second, we follow a similar methodology to analyze frontier AI’s impacts on defenses. We follow the convention to categorize defenses for systems and humans (Figure 4) and analyze each defense step through qualitative and quantitative analysis. Third, with a comprehensive understanding of frontier AI’s impact on attacks and defenses, we predict whether it will benefit attacks or defenses more. Finally, we discuss the marginal risks in hybrid systems (defined in Section 1) and potential mitigations.
We acknowledge that covering all risks and attack vectors is challenging and thus focus on critical attacks that have been impacted by AI. For example, we do not discuss attack vectors in cryptography, such as discovering vulnerabilities in cryptographic protocols. These attack vectors are currently beyond AI’s capabilities, though AI may be useful to such attacks later.
2.2 Qualitative & Quantitative Risk Analysis
For each attack and defense steps, our qualitative analysis exhaustively searches and systematically catalogs AI-facilitated techniques and instances in research papers and real-world deployment from the following four-tier impact: (1) no demonstrated effect, where we find no AI-related impact for this step; (2) demonstration in research papers, where we find only research papers but no real-world reports; (3) demonstration in real-world attacks, where we find research papers and a few early-stage real-world demonstrations; and (4) large-scale real-world deployments, where AI has been widely deployed in the real world.
Beyond qualitative analysis, we further study quantifying frontier AI’s impact in attacks and defenses. To this end, we collect existing benchmarks on AI-facilitated attacks and defenses and summarize their evaluation results. We also discuss the limitations of current benchmarks and how to comprehensively and accurately quantify frontier AI’s impact in the future. Building upon our qualitative and quantitative analysis, we explore the future impacts of frontier AI in cybersecurity and provide concrete recommendations for mitigating potential negative impacts. Table I presents a summary of the status quo.
TABLE I: Key summaries of the current status quo for each attack and defense step. The colors of the first column represent different levels of impact for the corresponding attack/defense step: Green–no demonstrated effect, Yellow–demonstrated in research papers, Orange–demonstrated in the real world, and Red–the large-scale real world.
  Attack steps Summary
@@ 2.2 Qualitative & Quantitative Risk Analysis
Step 1: Reconnaissance (Demonstrated in the real world) Qualitative: Both research and real-world attacks leverage LLMs in reconnaissance, including active scanning [270], victim information gathering [73, 200], and open-source database search [231].
Quantitative: Current benchmarks rely primarily on text-based assessments [274, 184, 286], while practical evaluations like AutoPenBench [101] show AI agents’ potential in reconnaissance but still limited.
@@
Step 2: Weaponization (Demonstrated in the real world) Qualitative: Research papers show LLMs can generate functional malware with high evasion rates [213] and AI agents aid in vulnerability identification and exploitation [186, 174, 90, 91], although large-scale real-world attacks remain limited to simpler attacks like DDoS [23, 281] and SQL injection [200, 119].
Quantitative: RedCode benchmark [111] reveals LLMs’ limitations in generating functional malware. Vulnerability exploitation benchmarks [286, 316, 251, 329] show Claude 3.5 Sonnet achieving cybersecurity practitioner-level skills but falling short of expert capabilities.
Qualitative: Initial access and exploitation (installation) are demonstrated in real world [96, 208, 142] while persistence remains at research level [245].
Step 3-5: Delivery & exploitation & installation (Demonstrated in the real world) Quantitative: Benchmark coverage and quality for these attack steps are very limited.
@@
Step 6: Command and control (Demonstrated in the real world) Qualitative: Most sub-categories have been demonstrated only in research papers, e.g., privilege escalation through exploit chain generation [68] and command and control through automated domain generation [245, 13], while credential access is used in real world [244, 148, 211].
Quantitative: Benchmark coverage and quality for this attack step are very limited.
@@
Step 7: Action on objectives (Large-scale real world impact) Real-world AI-enhanced attacks increase across various systems (Web, mobile, cloud) [23, 137, 63], with malicious purposes including malware deployment, business logic abuse, and credential theft, causing significant financial losses and data breaches [142, 191, 63].
@@
Attacks against humans (Large-scale real world impact) Frontier AI escalates attacks against humans, with studies showing increases in social engineering (135%) and voice phishing (260%) since ChatGPT’s adoption [67, 262, 131]. AI is misused for identity theft [192], deepfake-based fraud [49, 189], child exploitation [266], and psychological manipulation [238].
  Defense steps Summary
@@
Step 1: Proactive testing (Demonstrated in the real world) Qualitative: Research explores LLMs for proactive cybersecurity testing, including automated penetration testing [155, 73, 120] and vulnerability detection through code foundation models [80, 186] and AI-enhanced fuzzing [139, 310]. Real-world demonstrations exist but lack evidence of large-scale adoption.
Quantitative: AI penetration testing benchmarks remain nascent, with one end-to-end benchmark AutoPenBench [101]. Vulnerability detection benchmarks [80, 247, 199] show SOTA models achieve limited accuracy, but these benchmarks face challenges in data quality, limited code context, and insufficient test diversity.
Step 2: Attack detection (Demonstrated in the real world) Qualitative: Frontier AI addresses key limitations of traditional AI methods by eliminating manual feature engineering, reducing dependence on labeled datasets [282, 77], and improving generalizability to out-of-distribution data [117, 7]. Real-world applications exist in malware and network intrusion detection.
Quantitative: Many benchmarks exist for network intrusion and malware detection [252, 19, 315] but have data and label quality limitations. Foundation model-driven detection shows high accuracy [117, 178], calling for more challenging benchmarks.
Step 3: Triage & forensic (Demonstrated in research papers) Qualitative: Frontier AI usages are at an early stage, including LLMs for vulnerability analysis, improving symbolic execution with PoC generation [292, 126] and developing AI agents for root cause analysis [235]. Pre-trained foundation models demonstrate superior performance in binary analysis tasks [220, 305].
Quantitative: Recent benchmarks like CRUXEval [108] show promising results for PoC generation with GPT-4o achieving 75% pass@1 success rate, while no public benchmark exists for comprehensive AI-driven reverse engineering evaluation.
@@
Step 4-5: Remediation development & deployment (No demonstrated effect) Qualitative: Research demonstrates frontier AI can automatically generate security vulnerability patches [177, 159, 3, 34], though real-world applications like the SQLite3 Off-by-One bug fix [119] remain limited and no specific work exists on AI-assisted remediation deployment.
Quantitative: SOTA systems using Claude 3.5 Sonnet resolve about 50% of issues in SWE-bench-verified [147]. However, most SWE-bench issues are non-security bugs, revealing limitations in evaluating frontier AI’s defense capabilities.
@@
Defense for humans (Demonstrated in the real world) Frontier AI shows promise for enhancing defenses against human-targeted attacks, including social bot detection [94], fraud detection [326, 78], deepfake detection [229], and misinformation detection [14, 79, 180]. However, defensive techniques struggle to keep pace with sophisticated attacks [6], highlighting the need for research on adversarial dynamics between defensive and attack AI systems.
3 Frontier AI’s Impact on Cyberattacks

### v3

@@ Abstract
The impact of frontier AI in cybersecurity is rapidly increasing. In this paper, we comprehensively analyze this trend through three distinct lenses: a quantitative benchmark analysis, a literature review, and an expert survey. We find that while AI is already widely used in attacks, its application in defense remains limited, especially in remediation and deployment. Aligned with these analyses, experts expect AI to continue favoring attackers over defenders, though the gap will gradually narrow. These findings underscore the urgent need to mitigate frontier AI’s risks while closely monitoring emerging capabilities. We provide concrete calls-to-action regarding: the construction of new cybersecurity benchmarks, the development of AI agents for defense, the design of provably secure AI agents, the improvement of pre-deployment security testing and transparency, and the strengthening of user-oriented education and defenses. Our paper summary and blog are available at https://rdi.berkeley.edu/frontier-ai-impact-on-cybersecurity/.
**footnotetext: Co-first authors
@@ 1 Introduction
The security community is increasingly using frontier AI, [^1] We use frontier AI to represent foundation models like large language models (LLMs) and multi-modal models, as well as AI agents [66]—systems that combine foundation models with symbolic software components. e.g., using LLMs for fuzzing [156] and patching [276]. Researchers have also developed foundation models for security tasks such as analyzing networking traffic, binary code, and blockchain transactions, achieving better performance than traditional ML in threat detection [91, 77, 262, 179]. More recently, AI agents from the DARPA AIxCC competition [54] discovered and patched zero-day vulnerabilities [93]. Despite these defensive advances, frontier AI introduces significant risks, particularly its potential for misuse by attackers—ranging from the automation of cyber attacks [121, 236, 71, 223] to attacks against AI systems [237, 22], e.g, backdoors [164], jailbreaking [39, 40], and prompt injection [146].
Given this dual nature, a key question arises: Will frontier AI provide more benefits to attackers or defenders, and how might it reshape the cybersecurity landscape? While several recent papers examine LLMs in cybersecurity [73, 98, 81, 261, 193, 274, 280, 108, 205, 196], these studies do not address these questions: (1) How can we evaluate the current state of frontier AI in attacks and defenses? (2) How might the offense-defense balance change with frontier AI? and (3) What actions need to be taken to ensure AI benefits defense, increasing the security of systems? Answering these questions is especially critical to prepare for increasing access to frontier AI, amid the unprecedentedly rapid pace of AI evolvement (Figure 1[2, 202].
@@
Figure 1: Performance evolvement of frontier AI models over time. The coding-related tasks are emphasized in bold text.
Figure 2: A comprehensive categorization of attacks (red canvas) and defenses (blue canvas) with frontier AI’s qualitative impact. AI has a larger impact on final attack goals than on specific stages. The disparity exists because most reports disclose only final impacts rather than attack procedures. For precision, we include only reports with clear specifications for individual attack steps.
In this paper, we present a comprehensive analysis through benchmark analysis, a literature review, and an expert survey. First, we define the marginal risks of frontier AI to cybersecurity [122] and propose a risk taxonomy and risk assessment method (Section 2). We then apply this taxonomy to assess frontier AI’s current impacts on both attacks and defenses quantitatively and qualitatively (Sections 3 and 4). Our findings show that frontier AI has a much stronger influence on the attack side—particularly in human-targeted attacks—while its role in defense remains limited. In particular, remediation and deployment currently show no measurable impact. Through the expert survey, we also find that experts believe that frontier AI will benefit attackers more, even though the gap is expected to narrow over time; within the next two years, the balance is expected to favor attackers by $31.7\%,$ within five years by $12.8\%,$ and within ten years by just $0.46\%$ (Section 5). Our position aligns with this view, which we support through analysis of the equivalence class problem and the fundamental asymmetry between attackers and defenders (Section 6). Finally, we offer actionable recommendations to guide the community as follows (Section 7):
  • We recommend building security benchmarks with real-world systems and executable environments for agents.
  • We recommend improving traditional defenses with AI-based rule extraction and automation, as well as constructing AI agents with security-specific tools that automatically determine their workflows.
@@
  • We recommend AI developers conduct comprehensive pre-deployment security testing and enforce greater transparency in AI development.
  • We call for better education about AI-based attacks and security best practices, as well as developing human-oriented defenses.
In Appendix C, we list a set of open research questions about frontier AI and security.
@@ 2 Methodology
Based on the concept of marginal risk [122]—the new/unique risks beyond the existing risks—we define the marginal risks of frontier AI in cybersecurity as novel cyber threats either introduced or amplified by frontier AI through intentional misuse beyond traditional attack methods. For attacks against systems, we categorize attacks into seven steps based on the MITRE framework [161] and the cyber kill chain framework [151] (Figure 2); within each step, we define subcategories by activity and application. Similarly, we classify attacks against humans based on their goals and methods, as well as defenses for systems. We acknowledge that covering all risks and attack vectors is challenging; therefore, we limit our scope to cyberattacks against systems and humans (e.g, we do not discuss cryptographic attack vectors). With the taxonomy, we study frontier AI’s impact as follows:
Quantitative benchmarking. Because the evaluation of frontier AI’s cybersecurity capabilities is still at an early stage, available benchmarks remain limited. To collect benchmarks, we begin with several early and representative cybersecurity benchmarks (e.g., Cybench [272]), selected based on our domain expertise. We expand this set by adding other benchmarks cited within those works. We compile $34$ cybersecurity benchmarks and categorize them based on our taxonomy. We also identify the limitations of current benchmarks and propose calls-to-action to improve benchmarking in Section 7.
Literature review. We conduct a literature search using keywords related to frontier AI and cybersecurity in Google Scholar. The scope includes peer-reviewed papers, preprints, industry white papers, reports of real-world attack instances, technical blogs, and discussions from social and community platforms (e.g., X/Twitter) published between 2021 and Aug. 2025. This search produces an initial list of over 500 papers. In selecting articles, we prioritize conference publications, preprints from leading AI labs and major companies, and works with high citation impact. Five researchers participated in this process, resulting in a final list of $183$ papers. Through a qualitative analysis, we then map each attack and defense step to a four-tier impact: (1) no demonstrated effect, where we find no AI-related impact for this step; (2) demonstration in research papers, where we find only research papers but no real-world reports; (3) demonstration in real-world attacks, where we find research papers and a few early-stage real-world demonstrations; (4) large-scale real-world deployments, where AI has been widely deployed or had a noticeable real-world impact.
Expert survey. To assess how the AI/security community perceives the current and future impacts of frontier AI on cybersecurity, we conducted an expert survey targeting AI/security researchers and practitioners. The survey consists of questions about the current status quo of offensive and defensive use of frontier AI, predictions regarding timelines for automating cybersecurity tasks, and expectations about how frontier AI may shift the balance between attackers and defenders. We used snowball sampling to share the survey with experts; we emailed colleagues working at the intersection of AI and cybersecurity in academia and industry and asked them to circulate the survey within their professional networks. In total, $129$ experts accessed the survey, of whom $34$ completed it. Some questions were randomly assigned to respondents as part of the dynamic survey design, resulting in varying sample sizes across questions, with a maximum of $n=46.$ Survey questions, per-question sample sizes, and demographic characteristics (for respondents who provided them) are presented in Appendix B. We analyze and report experts’ forecasts of frontier AI’s future impacts and their views on the evolving offense-defense balance, including whether–and how–the perceived advantage for attackers may narrow over time.
3 Quantitative Analysis: Benchmarks
Table 1: Overview of existing benchmarks. Recon. - Reconnaissance, Mal. C - Malware creation, Vuln. E - Vulnerability exploitation, Ini. A & Del. - Initial access & delivery, Ex. & Inst. - Exploitation & installation, Per. & Ev. - Persistence & evasion, C2. & Act. - Command and control & action on objectives, Pen. Test - Penetration test, Vul. D - Vulnerability detection, Att. D - Attack detection, PoC & RT - PoC & Root cause, Rem. Dev. - Remediation development, Rem. Dep. - Remediation deployment. *Since Q&A benchmarks focus mainly on evaluating models’ cybersecurity knowledge, we excluded them from the table.
                                      Attack steps Quantity Notable examples Key limitations
1 Recon. 1 AutoPenBench [80] Limited method and scenario coverage; limited agent scaffolds
2
Mal. C 2 RedCode [88], CySecBench [235] Limited metric for malware creation,
Vul. E 7 CyberSecEval [236], BaxBench [232], BountyBench [270], CyberGym [242] limited program language coverage
3–5
Ini. A & Del. 2 AutoPenBench [80], SecCodePLT [260] Limited attack category and type coverage,
Ex. & Inst. 3 CyBench [272], SecCodePLT [260] limited benchmarks on real-world complex systems,
Per. & Ev. 0 None limited dynamic benchmarks with agent scaffold
6–7 C2. & Act. 1 SecCodePLT [260]
                                      Defense steps Quantity Notable examples Key limitations
1
Pen. test 1 AutoPenBench [80] Limited attacks, target systems, metrics,
Vul. D 12 PrimeVul [63], SVEN [100] FuzzBench [159], OSS-Fuzz [200] limited complexity and diversity, low-quality labels
2 Att. D 4 IDS2018 [204], Drebin [16], BODMAS [259] Low-quality data and labels, lack OOD/hidden tests
3
PoC & RT 2 CRUXEval [85], CyberGym [242] No benchmark for root cause
4 Rem. Dev. 5 SWE-bench [118], SecCodePLT [260], BountyBench [270] Limited benchmarks for vuln. in different languages
5 Rem. Dep. 0 None N/A
3.1 Offensive Benchmarks
As shown in Table 1, except for vulnerability exploitation in step 2, existing benchmarks are limited in risk coverage, metric design, and dynamic evolution. Table 5 in Appendix details SOTA AI’s performances on these attack benchmarks.
Attack Step 1: Reconnaissance. Early benchmarks focus on question answering [225, 147, 236, 124]. AutoPenBench [80] enables attack generation and execution but covers limited methods and scenarios. Results show that simple AI agents can handle reconnaissance like discovering potential victims on remote networks or conducting surveillance. However, agents cannot identify target-specific vulnerable services due to limited service information from standard tools (e.g., Nmap) and a lack of deeper interactions with target systems.
Attack Step 2: Weaponization. For malware creation, RedCode [88] tests LLMs’ capabilities in generating malware in Python and uses VirusTotal as the judge. CySecBench [235] tests the safety alignment of LLMs and shows that Claude 3.5 Sonnet is more resilient than GPT-4o and Gemini-1.5. Recent works design benchmarks for insecure code generation (e.g., CyberSecEval [236], SecCodePLT [260], BaxBench [232], SecRepoBench [62]). They craft specific coding tasks that may encounter known CWEs and evaluate whether LLMs generate vulnerable code when finishing these tasks, showing that both SOTA commercial and open-source models can generate vulnerable code, with stronger models being more vulnerable due to their enhanced task understanding and execution. BountyBench [270] and CVE-Bench [282] evaluate agent performance on vulnerability exploitation, showing that agents using Claude 3.7 Sonnet successfully exploit $67.5\%$ of tasks on BountyBench, while agents using GPT-4o achieve a $13\%$ success rate on CVE-Bench. CyberGym [242] evaluates agents’ capabilities in PoC generation on real-world software projects, identifying $15$ zero-day vulnerabilities. All existing benchmarks focus on C/C++ with limited coverage of other languages.
[…]
3.2 Defensive Benchmarks
Most benchmarks focus on vulnerability detection and attack detection, while other defense steps lack benchmarks. Table 6 in Appendix details AI’s performance on these benchmarks.
Defense Step 1: Proactive testing. Benchmarking penetration testing is related to attack benchmarking, which remains at an early stage with AutoPenBench [80]. Existing vulnerability detection benchmarks include static benchmarks with labeled functions [176, 63, 228, 35, 226, 234, 92, 172] and dynamic benchmarks with fuzzing instrumentation and execution environment [200, 159]. For static benchmarks, most existing benchmarks remain at the individual function level, e.g., PrimeVul [63] and SVEN [100]. These datasets have low-quality labels due to missing context (functions vulnerable only under conditions specified by other functions in the projection). SecCodePLT [260] includes a project-level vulnerability detection dataset for C/C++ and Java, showing the limited performance of existing models and agents in these languages. For dynamic benchmarks, OSS-Fuzz [200] is the largest continuous fuzzing platform for critical open-source projects. FuzzBench [159] is a smaller but less noisy benchmark widely used in research papers to test novel fuzzing methods, including AI-facilitated fuzzers [208, 275].
Defense Step 2: Attack detection. Many benchmarks exist for network intrusion detection and malware detection [204, 16, 259]. Recent research [15] highlights these benchmarks’ limitations in data quality (duplicated data, shortcuts) and label accuracy. Transformers show very high accuracy on these benchmarks [91, 140]. Besides attack detections, CTIBench [7] designs tasks for LLMs to pinpoint the specific CWEs and attack techniques for given attack instances.
Defense Step 3: Triage & forensic. CRUXEval [85] benchmarks PoC generation for vulnerable C/C++ codes with overflow bugs, showing that GPT-4o achieves 75% pass@1 success rate. CyberGym [242] offers more challenging tasks for agents, including PoC generation on real-world codebases with vulnerability descriptions, showing that OpenHands with Claude 4 Sonnet achieves a 17.9% success rate.
Defense Steps 4&5: Remediation development and deployment. Given that AI is not yet widely involved in these steps, there are not many benchmarks. The most popular benchmark is SWE-bench [118], which extracts issues from multiple GitHub Python projects. SOTA multi-agent systems with GPT-5 and Claude 4 Sonnet can resolve more than $70\%$ issues in the SWE-bench-verified benchmark. Recent works extend the SWE-bench to multi-modal settings [256] and the Java language [267]. SecCodePLT [260] (Python, C/C++, Java) and BountyBench [270] (C, Python, JavaScript/TypeScript) include benchmarks of security patches. Both benchmarks use unit tests to test functionality. SecCodePLT tests security via fuzzing through PoC, and BountyBench tests whether the patch prevents human-written exploits. Top agents (e.g., OpenAI Codex CLI o4-mini) can achieve high results on BountyBench (>90%), but remain ineffective on SecCodePLT (<30%).
Table 2: Key summaries of the qualitative literature review. The colors in the steps correspond to the same as in Figure 2. Literature supporting research impact and real-world impact is highlighted in blue and red, respectively.
  Attack steps Description Papers Impact Summary
@@ 3.2 Defensive Benchmarks
1: Reconnaissance Scan the env. or harvest info. to identify targets [221, 56, 189][160] LLMs in reconnaissance include active scanning, victim information gathering, and open-source database search.
@@
Step 2: Weaponization Create malware and exploits and encapsulate them as weapons [173, 205, 149, 135, 194, 63] [143, 236, 255, 242, 71, 270] [72, 175, 223, 272, 203] [20, 230, 160, 93, 115, 245] Research papers show LLMs can generate functional malware with high evasion rates and AI agents aid in vulnerability identification and exploitation, although large-scale real-world attacks remain limited to simpler attacks like DDoS and SQL injection.
Step 3: Delivery Transmit attacks through e-mail, USB, web [107, 238, 104, 181, 57, 53] [75, 169, 115, 160] Initial access are demonstrated in the real world, with AI assisting in password guessing and bypassing defenses such as 2FA.
Step 4: Exploitation Exploit vulnerabilities in the victim [52] AI can write simple exploits but cannot craft exploit chains.
Step 5: Installation Install the malware on the victim’s system [23][83, 199] AI is used for malware installation but not persistence in the real world.
@@
Step 6: Command & control Establish a remote command channel [95, 52, 227, 125, 96, 199, 11] [23, 43][119, 198, 171] Privilege escalation through exploit chain generation and command&control through automated domain generation are demonstrated in research. Credential access is used in the real world.
@@
Step 7: Action on objectives Achieve their ultimate objectives (e.g. data theft, sabotage) through the C&C channel [260, 250, 239, 229, 211] [20, 109, 230, 115, 75, 169, 48] [269, 78, 30, 152, 79] Real-world AI-enhanced attacks increase across various systems (Web, mobile, cloud), with malicious purposes including malware deployment, business logic abuse, and credential theft, causing significant financial losses and data breaches.
@@
Attacks against humans Exploit people often through deception, coercion, or manipulation [32, 70, 25, 101, 68, 14, 13, 82, 47] [182][36, 51, 213, 261, 197, 27] [153, 165, 210, 218, 106, 192, 18] Frontier AI significantly escalates attacks against humans (e.g., studies showing increases in social engineering and voice phishing since ChatGPT’s adoption).
  Defense steps Description Papers Impact Summary
@@
Step 1: Proactive testing Probe systems before an attack to uncover vulnerabilities [123, 94, 250, 224, 45, 105, 116] [56, 281, 252, 63, 4, 163, 194] [149, 135, 206, 183, 65, 220, 130] [205, 254, 168, 156, 58, 273, 275] [112, 136, 207, 155][269, 89] Research explores LLMs for proactive testing, including automated penetration testing and vulnerability detection and AI-enhanced fuzzing. Real-world demonstrations exist but lack evidence of large-scale adoption.
Step 2: Attack detection Identify malicious behavior as it happens and raise real-time alerts [240, 247, 34, 233, 127, 144, 140] [6, 91, 262, 77, 216, 266][33] LLMs remove manual feature engineering, reduce reliance on labeled datasets, and generalize better to out-of-distribution cases, with real-world use in malware and intrusion detection.
Step 3: Triage & forensic Reproduce the vulnerabilities and identify root causes [241, 99, 242, 166, 90, 162, 50] [191, 253, 184, 209, 42, 179, 248] [117, 251, 110][49] For PoC generation, recent works mainly explore LLM-assisted fuzzing. AI agents are developed for root cause analysis. Foundation models also show superior performance in binary analysis.
@@
Step 4: Rem. dev. Create patches or other countermeasures [133, 263, 126, 3, 157, 214, 41, 177] [276, 28, 139, 167, 265, 132][93] Research demonstrates AI can automatically generate security vulnerability patches, but real-world applications remain limited
Step 5: Rem. dep. Deploy those fixes in production NA No specific work exists on AI-assisted remediation deployment.
@@
Defense for humans Protect people from manipulation or exploitation in attacks [74, 268, 60, 31, 187, 5, 12] [61, 142, 277, 174, 141, 278, 87] [131, 158] Even though frontier AI shows promise for enhancing defenses against human-targeted attacks, defensive techniques struggle to keep pace with sophisticated attacks.
4 Qualitative Analysis: Literature Review
Table 2 summarizes the qualitative impact of frontier AI in attacks and defenses along with the categorized papers.